# Scalable piecewise smoothing with BART

### v3

@@ 3.1 The ridgeBART Prior
We specify a conjugate inverse gamma prior for $\sigma^{2}$ with the same default hyperparameters as Chipman et al., (2010). We specify the regression tree prior compositionally, by describing how to simulate a prior draw of a single regression tree. First, we draw the decision tree $\mathcal{T}$ from the branching process prior described in Section 2. Then, independently across leafs $\ell$ of $\mathcal{T},$ we draw each $\beta_{d,\ell}\sim\mathcal{N}\left(0,\tau^{2}\right)$ and each offset $b_{d,\ell}\sim\textrm{Uniform}\left(0,2\pi\right)$ if $\varphi$ is the cosine function of $b_{d,\ell}\sim\mathcal{N}\left(0,1\right)$ otherwise. Finally, we draw a leaf-specific scale parameter $\rho_{\ell}\sim\textrm{Gamma}\left(\nu/2,\nu\lambda/2\right)$ and then independently draw each $\omega_{d,\ell}\sim\mathcal{N}\left(0,\rho_{\ell}^{-1}\right).$
@@ 3.2 Sampling from the ridgeBART posterior
Unfortunately, it is not generally possible to obtain $p(\mathcal{T}|\bm{y},\mathcal{E}^{(-m)},\sigma^{2})$ in closed form, rendering the marginal decision tree and conditional jump updates impractical. Notice, however, that (i) we can easily marginalize out the outer weights $\bm{\beta}_{\ell}$’s from Equation 5 and (ii) that $\bm{\beta}_{\ell}$’s are conditionally independent and follow multivariate normal distributions given $\mathcal{T}$ and all the inner weights $\Xi=\{(\rho_{\ell},\bm{\omega}_{\ell},\bm{b}_{\ell})\}.$ Letting $P_{\ell}=\sigma^{-2}\Phi^{\top}_{\ell}\Phi_{\ell}+\tau^{-2}I_{n_{\ell}}$ and $\Theta_{\ell}=\sigma^{-2}\Phi_{\ell}^{\top}\bm{r}_{\ell},$ we compute
@@
In grow, prune, and change proposals, we draw the necessary new inner weights from the prior. That is, depending on the activation function, we draw the entries of $\bm{b}$ from either $\textrm{Uniform}\left(0,2\pi\right)$ or $\mathcal{N}\left(0,1\right).$ Then, we draw a new scale parameter $\rho$ before drawing the entries of $\bm{\omega}$ i.i.d. from a $\mathcal{N}\left(0,\rho^{-1}\right).$ The local nature of grow, prune, and change moves and the fact that we propose new inner weights from the prior facilitates considerable cancellation in the Metropolis-Hastings ratio.
Sampling $\sigma^{2}.$ After updating the regression trees in $\mathcal{E},$ we can update $\sigma^{2}$ with a conjugate normal-inverse gamma update.
Computational complexity. In each regression tree update, we must (i) compute and invert the matrix $P_{\ell},$ which takes $O(nD^{2}+D^{3})$ time; and (ii) evaluate the matrix-vector products $\Phi^{\top}\bm{r}_{\ell}$ and $P_{\ell}^{-1}\Theta_{\ell},$ which requires $O(n_{\ell}D+D^{2})$ time. When $D\ll n,$ the overall complexity of a ridgeBART regression tree update is effectively $O(n),$ exactly matching that of original BART.
@@ 4 Theoretical Results
In addition to its excellent empirical performance, the original BART model exhibits desirable theoretical properties: under a slight modification to the branching process prior for $\mathcal{T}$ and other mild assumptions, Ročková and Saha, (2019) showed that the BART posterior contracts at a nearly-minimax optimal rate when the regression function $f(\bm{x})$ belongs to the Hölder class $\mathcal{C}^{\alpha}([0,1]^{p}),$ for $10<\alpha<$ 1 (i.e., $f$ is rougher than Lipschitz). Jeong and Ročková, (2023) extended this result to the setting where $f(\bm{x})$ displays different levels of smoothness with respect to each covariate across different rectangular sub-regions of $[0,1]^{p}$ (i.e., when $f$ belongs to a piecewise anisotropic Hölder class; see Definition 2). Although their result is more general than Ročková and Saha, (2019), it is still limited by the assumption that $f$ is rougher than Lipschitz in each sub-region. In this section, we consider the setting where $\bm{x}=\bm{z}$ and show that for sufficiently smooth activation functions $\varphi,$ the ridgeBART posterior concentrates when $f(\bm{x})$ lies in a piecewise anisotropic Hölder smoothness class of arbitrary order.
The proof of our main result (Theorem 1) follows the general outline of the proofs in Jeong and Ročková, (2023). A key step involves approximating functions in a given class can be approximated arbitrarily well by a single deep regression tree, which can be further decomposed into a sum of shallower trees outputting linear combinations of sufficiently smooth ridge functions.
Setting and notation. Before formally defining our function space, we introduce some notation. Suppose that $\bm{\Psi}=[0,1]^{p}$ has been partitioned in $R$ non-overlapping rectangular boxes $\bm{\Psi}=\bigcup_{r=1}^{R}{\Psi_{r}}.$ We will allow the true regression function to display different degrees of smoothness within each region $\Psi_{r}.$ We denote the space of all continuous functions on $[0,1]^{p}$ by $C([0,1]^{p})$ and for any $h\in C([0,1]^{p}),$ $\bm{x}\in[0,1]^{p}$ and $1\leq i\leq p,$ let $h_{i}(\cdot|\bm{x})$ be the univariate slice function $y\mapsto h(x_{1},\ldots,x_{i-1},y,x_{i+1},\ldots,x_{p}).$ Additionally, for $\bm{n}=(n_{1},\dots,n_{p})$, let $D^{\bm{n}}h$ denote the mixed partial derivatives of order $(n_{1},\dots,n_{p})$ of $h$. Given covariate vectors $\bm{x}_{1},\ldots,\bm{x}_{n},$ we denote the empirical $L^{2}$ norm by $\lVert h\rVert_{n}^{2}=n^{-1}\sum_{i=1}^{n}h(\bm{x}_{i})^{2}$. Finally, given a vector $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{p})^{\top}\in\mathbb{R}^{p}_{+}$ let $\textrm{hm}\left(\bm{\alpha}\right)=\left(p^{-1}\sum_{j=1}^{p}{\alpha_{j}}\right)^{-1}$ be the harmonic mean of its entries.
@@
Bhattacharya et al., (2014) showed that posterior based on multi-bandwidth Gaussian Processes contract at near minimax-optimal rates when the true function lies in the anisotropic Hölder-class $\mathcal{H}^{\bm{\alpha},p}_{\lambda}(\Psi).$ We will instead consider the class of functions that, when restricted to each subregion $\Psi_{r},$ belong to an anisotropic Hölder class over each $\Psi_{r},$ but display different degrees of smoothness across each subregion. Definition 2 formally introduces our piecewise anisotropic Hölder class and is a natural extension of Jeong and Ročková, (2023)’s Definition 2.
@@
Note that although we allow $f\in\mathcal{H}^{\mathcal{A},p}_{\lambda}$ to exhibit different degrees of smoothness in each subregion $\Psi_{r},$ we will assume that $f$ is continuous across the boundaries of adjacent subregions.
The first key step in deriving ridgeBART’s posterior contraction rates over $\mathcal{H}^{\mathcal{A},p}_{\lambda}(\bm{\Psi})$ is to show that every function in the class is well-approximated by a linear combination of ridge functions.
@@
The proof of Lemma 1 is in Appendix C1 of the Supplementary Materials and follows the proof techniques in Yarotsky, (2017). Briefly, we begin by approximating the target function $f_{0}$ with polynomials in each box $\Psi_{r}.$ Then, using Lemma C1 in the Supplementary Materials and standard Hölder bounds, we show that each local polynomial approximates each $f_{0}|\Psi_{r}$ to within $D^{-\min_{j}{\alpha_{r,j}}/p}$ on each $\Psi_{r},$ where $D$ is the number of ridge units per leaf, as described in Equation 4. The total count of ridge expansions across $R$ subdomains is given by $D^{\star}\approx RD$. In simpler terms, our per-subdomain budget of ridge functions is $D$ and the total budget is $D^{\star}$. Using ideas from Barron, (1993) and Pinkus, (1997) we can approximate each polynomial path with a sum of ridge functions, while preserving the same approximation error order. Finally, we can “glue together” the local approximations into a single approximation by using a smooth partition of unity.
The next key result (Lemma C2 in the Supplementary Materials) is to show that for each $f\in\mathcal{H}^{\mathcal{A},p}_{\lambda}(\bm{\Psi}),$ there exists a sufficiently deep anisotropic k-d tree (Jeong and Ročková,, 2023, Definition 9) that outputs linear combinations of ridge functions that can approximate $f$ to within some error rate $\varepsilon_{n}$ that depends on $n,p,\lambda,$ and the harmonic mean of the vectors in $\mathcal{A}.$ This error rate turns out to be the posterior concentration rates.
To prove our main result, beyond assuming that the data arose from the model $y\sim\mathcal{N}\left(f_{0}(\bm{x}),\sigma_{0}^{2}\right)$ and that $f_{0}\in\mathcal{H}^{\mathcal{A},p}_{\lambda}(\bm{\Psi}),$ we make the following assumptions:
@@
    There are constants $a_{1}>0$ and $a_{2}>0$ such that $\min_{r,j}\alpha_{r,j}>a_{1}$ and $\lambda\lesssim n^{a_{2}}.$
@@
    The total number of ridge functions in $\hat{f}$, $D^{\star}\asymp\frac{n\epsilon_{n}^{2}}{\log n}.$
Assumption (A1) ensures the true function does not grow too quickly with $n$. This helps in prior mass arguments, where we need a region in parameter space that is large enough to contain $f_{0}$ but not so large as to hamper concentration. Similar boundedness assumptions appear in Ročková and Saha, (2019) and Jeong and Ročková, (2023). The assumption of $\min_{r,j}\alpha_{r,j}>a_{1}$ in (A2) ensures every subdomain $\Psi_{r}$ and coordinate $j$ has at least some minimum degree of smoothness. This excludes pathological cases with a near‐zero exponent that would severely degrade approximation rates. The Hölder coefficient $\lambda\lesssim n^{a_{2}}$ is a standard mild growth condition: the function can get slightly steeper with $n$, but not too rapidly. Assumption (A3) is the standard relationship balancing approximation and estimation errors in a sum‐of‐ridge or a neural network–like model.
@@
    In the branching process prior for $\mathcal{T},$ the probability that a node at depth $d$ is non-terminal is $\gamma^{d}$, with $21/n<\gamma<1/$ 2 .
@@
    The activation function $\varphi\in\mathcal{C}^{\bar{\alpha}^{-1}}(\bm{\Psi})$ and there exists a constant $C_{\varphi}$ not depending on $n$ or $p$ such that $\lVert\varphi\rVert_{\infty}<C_{\varphi}.$
Assumption (P1) is from Ročková and Saha, (2019), which is a standard way to control the distribution of tree shapes and sizes in the BART prior, ensuring enough splitting but not an explosion of large trees. Assumption (P2) controls the outer weights prior scale, preventing unbounded or negligible weights while Assumption (P3) is standard in Bayesian regression, ensuring good noise scale adaptivity. Assumption (P4) allows us to control the difference between evaluations of linear combinations of ridge functions with different inner and outer weights evaluated at different points.
Under these and three additional technical assumptions ((A4) – (A7) in Appendix C2 of the Supplementary Materials), the ridgeBART posterior places essentially all its mass in neighborhoods of $f_{0}$ and $\sigma_{0}^{2}$ whose empirical $L^{2}$ radius is $O(\epsilon_{n})$. Specifically, these additional assumptions ensure that there is regular-enough anisotropic k-d tree that can capture the behavior of $f_{0}$ in each sub-domain $\Psi_{r}$.
@@
The proof Theorem 1 is in Appendix C2 of the Supplementary Materials and follows the strategy in Jeong and Ročková, (2023) by verifying a Kullback-Leibler (KL), prior mass, and metric entropy conditions of Ghosal and van der Vaart, (2007, Theorem 4). Briefly, the KL condition ensures that there is a region in the parameter space where the KL divergence between the true data-generating model and the proposed model is small, allowing the posterior to concentrate around functions close to $f_{0}$. The prior mass condition guarantees that the prior assigns sufficient probability to neighborhoods of $f_{0}.$ The metric entropy condition limits the complexity of the function class by controlling the number of small neighborhoods (“sieves”) needed to cover the parameter space. Together, these conditions ensure the posterior concentrates around the true function.
Remark on harmonic mean equality.
In principle, the contraction can still be established without imposing $\textrm{hm}\left(\bm{\alpha}_{1}\right)=\cdots=\textrm{hm}\left(\bm{\alpha}_{R}\right).$ In that scenario, one must accommodate all subregions, making the global rate dependent on $\alpha_{\min}=\min_{1\leq r\leq R}\textrm{hm}\left(\bm{\alpha}_{r}\right).$ Concretely, under potentially different harmonic means $\textrm{hm}\left(\bm{\alpha}_{r}\right),$ the posterior converges at
$\epsilon_{n}~{}\asymp~{}(\lambda p)^{\tfrac{p}{(2\alpha_{\min}+p)}}\Bigl{(}\tfrac{R\,\log n}{n}\Bigr{)}^{\tfrac{\alpha_{\min}}{2\alpha_{\min}+p}}.$
However, any subregion with a smaller $\textrm{hm}\left(\bm{\alpha}_{r}\right)$ ultimately dictates the global speed, because it requires finer partitions to attain uniform precision, regardless of how smooth the other subregions may be. Moreover, balancing different depths or anisotropic splits across subregions introduces significant technical overhead. For these reasons, we impose the harmonic mean constraint so that each subregion shares the same “directional average” level of smoothness, thus yielding a simpler, unified approximation rate.
Remark on optimal recovery of isotropic Hölder functions.
When $R=1$ and all $\alpha_{1}=\cdots=\alpha_{p}=\alpha>0$, the space $\mathcal{H}^{\mathcal{A},p}_{\lambda}(\bm{\Psi})$ reduces to the usual isotropic Hölder class $\mathcal{C}^{\alpha}(\bm{\Psi}).$ Initial theoretical analyses of the BART posterior (e.g., Ročková and Saha,, 2019; Ročková and van der Pas,, 2019) showed that Bayesian ensembles of regression trees with constant leaf outputs could recover functions in $\mathcal{C}^{\alpha}(\bm{\Psi})$ at a near-optimal rate only for $0<\alpha\leq 1$ They further noted that the use of piecewise constant weak learners in the ensemble precludes near-optimal recovery of smoother functions. Our result, when specialized to the isotropic setting, goes further: by replacing constant outputs with linear combinations of sufficiently many sufficiently smooth ridge functions, ridgeBART can recover $\mathcal{C}^{\alpha}(\bm{\Psi})$ functions at a near-optimal rate for arbitrary $\alpha$.

### v4

@@ 3.1 The ridgeBART Prior
We specify a conjugate inverse gamma prior for $\sigma^{2}$ with the same default hyperparameters as Chipman et al., (2010). We specify the regression tree prior compositionally, by describing how to simulate a prior draw of a single regression tree. First, we draw the decision tree $\mathcal{T}$ from the branching process prior described in Section 2. Then, independently across leafs $\ell$ of $\mathcal{T},$ we draw each $\beta_{d,\ell}\sim\mathcal{N}\left(0,\tau^{2}\right)$ and each offset $b_{d,\ell}\sim\textrm{Uniform}\left(0,2\pi\right)$ if $\varphi$ is the cosine function and $b_{d,\ell}\sim\mathcal{N}\left(0,1\right)$ otherwise. Finally, we draw a leaf-specific scale parameter $\rho_{\ell}\sim\textrm{Gamma}\left(\nu/2,\nu\lambda/2\right)$ and then independently draw each $\omega_{d,\ell}\sim\mathcal{N}\left(0,\rho_{\ell}^{-1}\right).$
@@ 3.2 Sampling from the ridgeBART posterior
Unfortunately, it is not generally possible to obtain $p(\mathcal{T}|\bm{y},\mathcal{E}^{(-m)},\sigma^{2})$ in closed form, rendering the marginal decision tree and conditional jump updates impractical. Notice, however, that (i) we can easily marginalize out the outer weights $\bm{\beta}_{\ell}$’s from Equation 5 and (ii) that the $\bm{\beta}_{\ell}$’s are conditionally independent and follow multivariate normal distributions given $\mathcal{T}$ and all the inner weights $\Xi=\{(\rho_{\ell},\bm{\omega}_{\ell},\bm{b}_{\ell})\}.$ Letting $P_{\ell}=\sigma^{-2}\Phi^{\top}_{\ell}\Phi_{\ell}+\tau^{-2}I_{n_{\ell}}$ and $\Theta_{\ell}=\sigma^{-2}\Phi_{\ell}^{\top}\bm{r}_{\ell},$ we compute
@@
In grow, prune, and change proposals, we draw new inner weights from the prior. While it is tempting to propose $\Xi$ in a more data-informed fashion (e.g., by searching for an optimal set of inner weights), doing so yields proposals with exceedingly small acceptance probabilities, which in turn prevents the Markov chain from moving (see Deshpande,, 2024, Appendix B2 for further details). The local nature of grow, prune, and change moves and the fact that we propose new inner weights from the prior facilitates considerable cancellation in the Metropolis-Hastings ratio.
Sampling $\sigma^{2}.$ After updating $\mathcal{E},$ we draw a new $\sigma^{2}$ from a conditionally conjugate inverse gamma distribution.
Computational complexity. In each regression tree update, we must (i) compute and invert the matrix $P_{\ell},$ which takes $O(n_{\ell}D^{2}+D^{3})$ time; and (ii) evaluate the matrix-vector products $\Phi^{\top}\bm{r}_{\ell}$ and $P_{\ell}^{-1}\Theta_{\ell},$ which requires $O(n_{\ell}D+D^{2})$ time. Thus, updating a single regression tree with $L$ leaves takes $O(n(D^{2}+D)+L(D^{3}+D^{2}))$ time. In practice, we recommend setting $D=1$ and typically observe $L\ll n.$ Consequently, the typical ridgeBART tree update’s complexity is effectively $O(n),$ which is exactly the same as the original BART.
By updating the inner weights $\Xi$ alongside the decision tree structure, we can fit much more expressive piecewise smooth models using a sampler that is nearly identical to that of the original BART model, which could only represent piecewise constant functions. In other words, we have vastly expanded the representational flexibility of BART without introducing substantial computational burden.
@@ 4 Theoretical Results
In addition to its excellent empirical performance, the original BART model exhibits desirable theoretical properties: under a slight modification to the branching process prior for $\mathcal{T}$ and other mild assumptions, Ročková and Saha, (2019) showed that the BART posterior contracts at a nearly-minimax optimal rate when the regression function $f(\bm{x})$ belongs to the Hölder class $\mathcal{C}^{\alpha}([0,1]^{p}),$ for $0<\alpha<1$ (i.e., $f$ is rougher than Lipschitz). Jeong and Ročková, (2023) extended this result to the setting where $f(\bm{x})$ displays different levels of smoothness with respect to each covariate across different rectangular sub-regions of $0,1]^{p}$ (i.e., when $f$ belongs to a piecewise anisotropic Hölder class; see Definition 2). Their more general result is nevertheless limited by the assumption that $f$ is rougher than Lipschitz in each sub-region. In this section, we consider the setting where $\bm{x}=\bm{z}$ and show that for sufficiently smooth activation functions $\varphi,$ the ridgeBART posterior concentrates when $f(\bm{x})$ lies in a piecewise anisotropic Hölder smoothness class of arbitrary order.
Setting and notation. Before formally defining our function space, we introduce some notation. Suppose that $\bm{\Psi}=[0,1]^{p}$ has been partitioned in $R$ non-overlapping rectangular boxes $\bm{\Psi}=\bigcup_{r=1}^{R}{\Psi_{r}}.$ We will allow the true regression function to display different degrees of smoothness within each region $\Psi_{r}.$ We denote the space of all continuous functions on $0,1]^{p}$ by $C([0,1]^{p})$ and for any $h\in C([0,1]^{p}),$ $\bm{x}\in[0,1]^{p}$ and $1\leq i\leq p,$ let $h_{i}(\cdot|\bm{x})$ be the univariate slice function $y\mapsto h(x_{1},\ldots,x_{i-1},y,x_{i+1},\ldots,x_{p}).$ Additionally, for $\bm{n}=(n_{1},\dots,n_{p})$, let $D^{\bm{n}}h$ denote the mixed partial derivatives of order $(n_{1},\dots,n_{p})$ of $h$. Given covariate vectors $\bm{x}_{1},\ldots,\bm{x}_{n},$ we denote the empirical $L^{2}$ norm by $\lVert h\rVert_{n}^{2}=n^{-1}\sum_{i=1}^{n}h(\bm{x}_{i})^{2}$. Finally, given a vector $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{p})^{\top}\in\mathbb{R}^{p}_{+}$ let $\textrm{hm}\left(\bm{\alpha}\right)=p\left(\sum_{j=1}^{p}\alpha_{j}^{-1}\right)^{-1}$ be the harmonic mean of its entries.
@@
Bhattacharya et al., (2014) showed that posteriors based on multi-bandwidth Gaussian Processes contract at near minimax-optimal rates when the true function lies in the anisotropic Hölder-class $\mathcal{H}^{\bm{\alpha},p}_{\lambda}(\Psi).$ We will instead consider the class of functions that, when restricted to each subregion $\Psi_{r},$ belong to an anisotropic Hölder class over each $\Psi_{r},$ but display different degrees of smoothness across each subregion. Definition 2 formally introduces our piecewise anisotropic Hölder class and is a natural extension of Jeong and Ročková, (2023)’s Definition 2. Although we allow $f\in\mathcal{H}^{\mathcal{A},p}_{\lambda}$ to exhibit different degrees of smoothness in each subregion $\Psi_{r},$ we will assume that $f$ is continuous across the boundaries of adjacent subregions.
@@
The first key step in deriving ridgeBART’s posterior contraction rate over $\mathcal{H}^{\mathcal{A},p}_{\lambda}(\bm{\Psi})$ is to show that every function in the class is well-approximated by a linear combination of ridge functions.
@@
The proof of Lemma 1 is in Section C.1 and follows the proof techniques in Leshno et al., (1993), Barron, (1993), and DeVore and Lorentz, (1993, Chapters 7 & 8). Briefly, we begin by approximating the target function $f_{0}$ with polynomials in each box $\Psi_{r}.$ Then, using Lemma C1 and standard Hölder bounds, we show that each local polynomial approximates each $f_{0}|\Psi_{r}$ to within $D^{-\min_{j}{\alpha_{r,j}}/p}$ on each $\Psi_{r},$ where $D$ is the number of ridge units per leaf, as described in Equation 4. The total count of ridge expansions across $R$ subdomains is given by $D^{\star}\approx RD$. In simpler terms, our per-subdomain budget of ridge functions is $D$ and the total budget is $D^{\star}$. Using ideas from Leshno et al., (1993) and Pinkus, (1997) we can approximate each polynomial path with a sum of ridge functions, while preserving the same approximation error order. Finally, we can “glue together” the local approximations into a single approximation by using a smooth partition of unity.
The next key result (Lemma C2) shows that for each $f\in\mathcal{H}^{\mathcal{A},p}_{\lambda}(\bm{\Psi}),$ there is a sufficiently deep anisotropic k-d tree (Jeong and Ročková,, 2023, Definition 9) outputting linear combinations of ridge functions that approximates $f$ to within some error rate $\varepsilon_{n}.$ This error rate, which depends on $n,p,\lambda,$ and the harmonic mean of the vectors in $\mathcal{A},$ turns out to be the posterior concentration rate.
To prove our main result, beyond assuming that the data arose from the model $y\sim\mathcal{N}\left(f_{0}(\bm{x}),\sigma_{0}^{2}\right)$ with $f_{0}\in\mathcal{H}^{\mathcal{A},p}_{\lambda}(\bm{\Psi}),$ we make the following assumptions:
@@
    The Hölder coefficient $\lambda$ has at most polynomial growth in $n$, i.e., $\lambda\lesssim n^{a_{2}}$ for some $a_{2}>0$.
@@
    The number of ridge functions in $\hat{f}$, $D^{\star}\asymp R\cdot\epsilon_{n}^{-p/\bar{\alpha}}$.
Assumption (A1) mildly restricts the supremum norm of $f_{0}$, which is standard in high-dimensional Bayesian regression analyses, to guarantee sufficient prior concentration under the normal prior (Jeong and Ročková,, 2023, c.f. Assumption A3). Similar boundedness assumptions appear in Ročková and Saha, (2019) and Jeong and Ročková, (2023). Assumption (A2) with the Hölder coefficient $\lambda\lesssim n^{a_{2}}$ is a standard mild growth condition: the function can get slightly steeper with $n$, but not too rapidly. Assumption (A3) is the standard relationship balancing approximation and estimation errors in a sum‐of‐ridge or a neural network–like model.
@@
    In the branching process prior for $\mathcal{T},$ the probability that a node at depth $d$ is non-terminal is $\gamma^{d}$, with $1/n<\gamma<1/2.$
@@
    The activation function $\varphi\in\mathcal{C}^{k}(\mathbb{R})$ for some integer $k\geq 1;$ is not a polynomial; and its Lipschitz constant on any compact set is bounded by $L_{\varphi}$.
Assumption (P1) is from Ročková and Saha, (2019), which is a standard way to control the distribution of tree shapes and sizes in the BART prior. Assumption (P2) controls the outer weights prior scale, preventing unbounded or negligible weights while Assumption (P3) is standard in Bayesian regression, ensuring good noise scale adaptivity. Assumption (P4) is central to the model’s expressive power. By requiring $\varphi$ to be sufficiently smooth, we ensure the ridge function building blocks can themselves represent smooth functions. The requirement that it is non-polynomial is also critical (stated in Leshno et al., (1993, Theorem 1)), as a polynomial activation function would limit the model to only generating other polynomials, thereby losing the universal approximation capability needed to represent arbitrary functions. This is the key mechanism that allows ridgeBART to approximate target functions of arbitrary smoothness $(\alpha>0),$ overcoming the limitations of standard BART.
Under these and five additional technical assumptions ((A4) – (A8) in Section C.2, the ridgeBART posterior places essentially all its mass in neighborhoods of $f_{0}$ and $\sigma_{0}^{2}$ whose empirical $L^{2}$ radius is $O(\epsilon_{n})$. These additional assumptions ensure that there is regular-enough anisotropic k-d tree that can capture the behavior of $f_{0}$ in each sub-domain $\Psi_{r}$.
@@
The proof Theorem 1 is in Section C.2 and follows the strategy in Jeong and Ročková, (2023) by verifying a Kullback-Leibler (KL), prior mass, and metric entropy conditions of Ghosal and van der Vaart, (2007, Theorem 4). We note that the contraction rate $\epsilon_{n}$ can, in principle, be derived without the assuming $\textrm{hm}\left(\bm{\alpha}_{1}\right)=\cdots=\textrm{hm}\left(\bm{\alpha}_{R}\right).$ In that more general scenario, the global posterior contraction rate is dictated by the function restricted to the sub-region over which it is the least smooth.
When $R=1$ and all $\alpha_{1}=\cdots=\alpha_{p}=\alpha>0$, the space $\mathcal{H}^{\mathcal{A},p}_{\lambda}(\bm{\Psi})$ reduces to the usual isotropic Hölder class $\mathcal{C}^{\alpha}(\bm{\Psi}).$ Initial theoretical analyses of the BART posterior (e.g., Ročková and Saha,, 2019; Ročková and van der Pas,, 2019) showed that Bayesian ensembles of regression trees with constant leaf outputs could recover functions in $\mathcal{C}^{\alpha}(\bm{\Psi})$ at a near-optimal rate only for $0<\alpha\leq 1.$ They further noted that the use of piecewise constant weak learners in the ensemble precludes near-optimal recovery of smoother functions. By using ensembles of soft decision trees, which output smooth combinations of all leaf outputs, Linero and Yang, (2018)’s SoftBART can recover $\mathcal{C}^{\alpha}(\bm{\Psi})$ functions for arbitrary $\alpha\geq 0$ Our result, when specialized to the isotropic setting, matches their result: by replacing constant outputs of hard decision trees with linear combinations of sufficiently many sufficiently smooth ridge functions, ridgeBART can recover $\mathcal{C}^{\alpha}(\bm{\Psi})$ functions at a near-optimal rate for any $\alpha\geq 0.$